### v1

# Spectrum and Resonances of Stark Hamiltonians with $\delta$–Shell Potentials

Abstract
We study the Stark Hamiltonian with a $\delta$–interaction supported on a hypersphere, with boundary strength $\alpha\in L^{\infty}(S_{a})$. Using the framework of quasi boundary triples, we construct self–adjoint realizations and derive a Krein–type resolvent formula. For $F\neq 0$ we prove $\sigma(H_{F,\alpha})=\sigma_{\mathrm{ac}}(H_{F,\alpha})=\mathbb{R}$ by transferring a strict Mourre estimate from the free Stark operator and by the compactness of the boundary coupling. Resonances are defined by complex distortion in the field direction and characterized as zeros of a boundary regularized Fredholm determinant. In $d=3$ and small $F$, bound states at $F=0$ continue to resonances with widths of order $\exp(-c/F)$, governed by an Agmon–type distance. A brief comparison with the free Stark case is also given.
1 Introduction
Two basic models in spectral and scattering theory motivate this work. The free Stark Hamiltonian
$H_{F,0}=-\Delta+Fx_{1},\qquad F\in\mathbb{R},$
acting on $L^{2}(\mathbb{R}^{d})$ for $d\geq 2$, is self–adjoint with purely absolutely continuous spectrum and satisfies the limiting absorption principle; see Herbst [12, 13] and Bentosela–Carmona–Duclos–Simon–Souillard–Weder [4]. It is also a standard model of delocalization under a constant electric field.
The penetrable–wall (or $\delta$–shell) Hamiltonian
$H=-\Delta+\alpha(\omega)\,\delta(|x|-a),$
is defined through the closed form
$h[u,v]=(\nabla u,\nabla v)_{L^{2}(\mathbb{R}^{d})}+(A_{\alpha}\gamma u,\gamma v)_{L^{2}(S_{a})},\qquad D[h]=H^{1}(\mathbb{R}^{d}),$
where $\gamma:H^{1}(\mathbb{R}^{d})\to L^{2}(S_{a})$ is the trace on $S_{a}=\{x:|x|=a\}$ and $A_{\alpha}$ denotes multiplication by $\alpha\in L^{\infty}(S_{a};\mathbb{R})$. We use the standard Sobolev spaces $H^{s}(\mathbb{R}^{d})$ ($s=1,2$), and on $S_{a}$ the spaces $H^{\pm 1/2}(S_{a})$; here $H^{1/2}(S_{a})=\mathrm{ran}\,\gamma$ and $H^{-1/2}(S_{a})$ is its $L^{2}(S_{a})$–dual. These conventions will be used without further comment. In the case $d=3$, the spectrum of $H$ equals $0,\infty)$ together with finitely many negative eigenvalues; these negative eigenvalues correspond to bound states created by the wall [18, 14].
The two models show opposite tendencies: a constant field favors delocalization and destroys bound states, whereas a $\delta$–shell creates localized states below the continuum. Their combination
$H_{F,\alpha}=H_{F,0}+A_{\alpha}\,\delta(|x|-a)$
naturally raises the question of how the balance between delocalization by the field and localization at the wall appears in the spectral and resonance structure. Heuristically, bound states for $F=0$ should dissolve into the continuum for $F\neq 0$ and persist only as resonances. For the free Stark Hamiltonian such resonances are known only in an abstract sense with exponentially small widths, while in the present setting they can be traced explicitly as continuations of bound states.
The aim of this paper is to give a systematic analysis of $H_{F,\alpha}$ by combining the boundary operator (quasi boundary triple) framework with spectral deformation. Our main results are as follows.
  • (1)
    Self–adjoint realization and Krein–type resolvent formula. A self–adjoint realization is constructed within the quasi boundary triple framework, and a Krein–type resolvent formula is derived in terms of a Poisson operator and a boundary Weyl function associated with the free Stark resolvent, with the boundary coupling entering as the multiplier $A_{\alpha}$ on $L^{2}(S_{a})$.
  • (2)
    Spectrum for $F\neq 0$. We prove $\sigma(H_{F,\alpha})=\sigma_{\mathrm{ac}}(H_{F,\alpha})=\mathbb{R}$ by transferring a strict Mourre estimate from $H_{F,0}$ and using compactness of the boundary coupling, so that embedded eigenvalues and singular continuous spectrum are excluded.
  • (3)
    Resonances. Resonances are defined by complex distortion in the field direction and are characterized as zeros of a boundary regularized Fredholm determinant, which gives an effective boundary criterion for detection.
  • (4)
    Small–field asymptotics in $d=3$. For small $F>0$, bound states at $F=0$ continue to resonances whose widths are exponentially small in $1/F$, governed by an Agmon–type distance.
2 Boundary triples and self–adjoint extensions
To analyze the self–adjoint extensions of the symmetric operator obtained by restricting $H_{F,0}$ outside the interaction surface $S_{a}$, we employ the framework of (quasi) boundary triples. This abstract machinery has two main advantages. First, it provides a convenient way to encode the boundary conditions in terms of boundary operators $\Gamma_{0},\Gamma_{1}$ that satisfy an abstract Green’s identity. Second, it allows one to parametrize all self–adjoint extensions and to derive Krein–type resolvent formulas in a transparent manner. In particular, the choice of the boundary condition $\Gamma_{0}=A_{\alpha}^{b}\Gamma_{1}$ will lead to the penetrable–wall Hamiltonian $H_{F,\alpha}$, where $A_{\alpha}^{b}:H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ denotes multiplication by $\alpha$ followed by the canonical embedding $L^{2}(S_{a})\hookrightarrow H^{-1/2}(S_{a})$. Passing to the $L^{2}(S_{a})$ pivot we set $A_{\alpha}^{\flat}:=\Lambda^{-1}\widehat{A}_{\alpha}\Lambda$.
We briefly recall the notion of a quasi boundary triple. Let $T$ be a densely defined symmetric operator with adjoint $T^{*}$. A triple $(\mathcal{G},\Gamma_{0},\Gamma_{1})$, where $\mathcal{G}$ is a Hilbert space and $\Gamma_{0},\Gamma_{1}:D(T^{*})\to\mathcal{G}$ are linear mappings, is called a quasi boundary triple for $T^{*}$ if the following hold:
  • the Green’s identity
    $(T^{*}u,v)-(u,T^{*}v)=(\Gamma_{1}u,\Gamma_{0}v)_{\mathcal{G}}-(\Gamma_{0}u,\Gamma_{1}v)_{\mathcal{G}},\qquad u,v\in D(T^{*}),$
    is satisfied,
  • the operator $T^{*}\upharpoonright\ker\Gamma_{0}$ is self–adjoint.
Here and in the sequel, $D(A)$ denotes the domain of the operator $A$. This concept provides a functional framework for parametrizing self–adjoint extensions and for deriving Krein–type resolvent formulas (see e.g. Behrndt–Hassi–de Snoo [3]).
Let
$H_{F,0}=-\Delta+Fx_{1}\ \ \text{on }L^{2}(\mathbb{R}^{d}),\ \ D(H_{F,0})=H^{2}(\mathbb{R}^{d}),\ d\geq 2.$
Restrict $H_{F,0}$ to $C_{0}^{\infty}(\mathbb{R}^{d}\setminus S_{a})$ and denote
$T:=H_{F,0}\upharpoonright C_{0}^{\infty}(\mathbb{R}^{d}\setminus S_{a}).$
Then $T$ is densely defined and symmetric.
Lemma 1 (Domain of the adjoint and trace data).
The adjoint $T^{*}$ has domain
$D(T^{*})=\{u\in L^{2}(\mathbb{R}^{d}):\ u|_{\{|x|<a\}},\ u|_{\{|x|>a\}}\in H^{2}_{\mathrm{loc}},\ H_{F,0}u\in L^{2}(\mathbb{R}^{d})\}.$
For each $u\in D(T^{*})$, the one–sided traces $u^{\pm}|_{S_{a}}$ belong to $H^{1/2}(S_{a})$, the one–sided normal derivatives $(\partial_{r}u)^{\pm}|_{S_{a}}$ belong to $H^{-1/2}(S_{a})$, and the jump $\partial_{r}u]:=(\partial_{r}u)^{+}-(\partial_{r}u)^{-}\in H^{-1/2}(S_{a})$ is well defined. Here $\partial_{r}$ denotes the outward normal derivative, and + (resp. –) refers to the exterior (resp. interior) side.
Proof.
Let $u\in D(T^{*})$. By definition there exists $f\in L^{2}$ such that $(Tv,u)=(v,f)$ for all $v\in C_{0}^{\infty}(\mathbb{R}^{d}\setminus S_{a})$. This means $H_{F,0}u=f$ in the distribution sense on $\mathbb{R}^{d}\setminus S_{a}$. Since $\Delta+Fx_{1}$ is strongly elliptic with bounded coefficients on the interior and exterior domains
$D^{-}:=\{\,x\in\mathbb{R}^{d}:|x|<a\,\},\qquad D^{+}:=\{\,x\in\mathbb{R}^{d}:|x|>a\,\},$
elliptic regularity yields $u|_{D^{\pm}}\in H^{2}_{\mathrm{loc}}(D^{\pm})$ whenever $u\in L^{2}$ and $H_{F,0}u\in L^{2}$. On Lipschitz boundaries, the trace map $H^{1}(\Omega)\to H^{1/2}(\partial\Omega)$ is continuous, and for $H^{2}_{\mathrm{loc}}$ solutions the conormal derivative belongs to $H^{-1/2}(\partial\Omega)$. Hence $u^{\pm}|_{S_{a}}\in H^{1/2}(S_{a})$ and $(\partial_{r}u)^{\pm}|_{S_{a}}\in H^{-1/2}(S_{a})$; the jump $\partial_{r}u]$ is then defined as the difference of these one–sided traces. $\Box$
We use the one–sided traces $u^{\pm}|_{S_{a}}\in H^{1/2}(S_{a})$ and the one–sided normal derivatives $(\partial_{r}u)^{\pm}|_{S_{a}}\in H^{-1/2}(S_{a})$ of $u\in D(T^{*})$. Set the average trace
$\gamma_{\mathrm{av}}u:=\tfrac{1}{2}\big{(}u^{+}|_{S_{a}}+u^{-}|_{S_{a}}\big{)}\in H^{1/2}(S_{a}),$
and the jump of the normal derivative
$\partial_{r}u]:=(\partial_{r}u)^{+}-(\partial_{r}u)^{-}\in H^{-1/2}(S_{a}).$
We adopt the following convention for the boundary operators:
$\Gamma_{0}u:=[\partial_{r}u]\in H^{-1/2}(S_{a}),\qquad\Gamma_{1}u:=\gamma_{\mathrm{av}}u\in H^{1/2}(S_{a}).$
Here $\Gamma_{0}$ corresponds to the normal jump across $S_{a}$, while $\Gamma_{1}$ is the average trace.
Lemma 2 (Green’s identity and quasi boundary triple).
For all $u,v\in D(T^{*})$,
$(T^{*}u,v)-(u,T^{*}v)=\langle\Gamma_{1}u,\Gamma_{0}v\rangle_{S_{a}}-\langle\Gamma_{0}u,\Gamma_{1}v\rangle_{S_{a}},$ (2.1)
where $\langle\cdot,\cdot\rangle_{S_{a}}$ denotes the dual pairing $H^{1/2}(S_{a})\times H^{-1/2}(S_{a})\to\mathbb{C}$. In particular, $(H^{-1/2}(S_{a}),\Gamma_{0},\Gamma_{1})$ forms a quasi boundary triple for $T^{*}$ in the sense of Behrndt–Hassi–de Snoo [3].
Proof.
Fix $\varepsilon>0$ small. Choose $\chi_{\pm}\in C^{\infty}(\mathbb{R}^{d})$ with $\chi_{+}+\chi_{-}=1$, $\chi_{+}=1$ on $|x|\geq a+\varepsilon$ and $\chi_{-}=1$ on $|x|\leq a-\varepsilon$. These cutoffs separate $D^{\pm}$ and allow us to treat the two sides of $S_{a}$ independently. Approximate $u,v\in D(T^{*})$ by sequences in $C_{0}^{\infty}(D^{+})\oplus C_{0}^{\infty}(D^{-})$ in the graph norms. Integration by parts on $D^{\pm}$ gives
$(T^{*}u,v)-(u,T^{*}v)$ $=$ $\langle(\partial_{r}u)^{+},v^{+}|_{S_{a}}\rangle_{S_{a}}-\langle u^{+}|_{S_{a}},(\partial_{r}v)^{+}\rangle_{S_{a}}$
$-\langle(\partial_{r}u)^{-},v^{-}|_{S_{a}}\rangle_{S_{a}}+\langle u^{-}|_{S_{a}},(\partial_{r}v)^{-}\rangle_{S_{a}}.$
Rearranging the terms, one obtains
$\langle(\partial_{r}u)^{+}-(\partial_{r}u)^{-},\tfrac{1}{2}(v^{+}|_{S_{a}}+v^{-}|_{S_{a}})\rangle_{S_{a}}-\langle\tfrac{1}{2}(u^{+}|_{S_{a}}+u^{-}|_{S_{a}}),(\partial_{r}v)^{+}-(\partial_{r}v)^{-}\rangle_{S_{a}}.$
Thus the expression matches precisely the right-hand side of (2.1), because $\Gamma_{0}$ is defined as the jump of the normal derivatives and $\Gamma_{1}$ as the average of the two traces. Since the computation was first carried out for smooth approximants, taking the limit in the graph norm yields the identity for all $u,v\in D(T^{*})$, and the proof is complete. $\Box$
Lemma 3 (Interior–exterior Dirichlet problems for nonreal energy).
Let $z\in\mathbb{C}\setminus\mathbb{R}$. For each $\phi\in H^{1/2}(S_{a})$ there exist unique $u^{\pm}\in H^{2}_{\mathrm{loc}}(D^{\pm})\cap L^{2}(D^{\pm})$ solving $(H_{F,0}-z)u^{\pm}=0$ in $D^{\pm}$ with trace $u^{\pm}|_{S_{a}}=\phi$. The solutions depend continuously on $\phi$, and interior regularity holds away from $S_{a}$.
Proof.
Consider first the interior region $D^{-}$. Let $E_{\mathrm{tr}}:H^{1/2}(S_{a})\to H^{1}(D^{-})$ be a fixed right inverse of the trace operator. We then decompose the solution as
$u^{-}=E_{\mathrm{tr}}\phi+w,\qquad w\in H^{1}_{0}(D^{-}).$
Substituting this form into the weak formulation and testing against $v\in H^{1}_{0}(D^{-})$ yields
$\int_{D^{-}}\nabla w\cdot\nabla\overline{v}\,dx+\int_{D^{-}}(Fx_{1}-z)w\,\overline{v}\,dx$
$=$ $-\int_{D^{-}}\nabla(E_{\mathrm{tr}}\phi)\cdot\nabla\overline{v}\,dx-\int_{D^{-}}(Fx_{1}-z)E_{\mathrm{tr}}\phi\,\overline{v}\,dx.$
The sesquilinear form
$q(w,v)=(\nabla w,\nabla v)_{D^{-}}+((Fx_{1}-\text{{\rm Re}}z)w,v)_{D^{-}}-i\,\text{{\rm Im}}~z\,(w,v)_{D^{-}}$
is continuous on $H^{1}_{0}(D^{-})$. Moreover, since $x_{1}$ is bounded on $D^{-}$, there exists $\lambda>0$ such that
$\text{{\rm Re}}\,q(w,w)+\lambda\|w\|_{L^{2}(D^{-})}^{2}\geq\tfrac{1}{2}\|\nabla w\|_{L^{2}(D^{-})}^{2}+\tfrac{1}{2}\lambda\|w\|_{L^{2}(D^{-})}^{2}.$
This shows that $q$ is coercive on $H^{1}_{0}(D^{-})$. Hence, by the Lax–Milgram theorem, there exists a unique $w\in H^{1}_{0}(D^{-})$ solving the weak problem, and the solution depends continuously on $\phi$. Elliptic interior regularity implies $u^{-}\in H^{2}_{\mathrm{loc}}(D^{-})$.
We consider next the exterior region $D^{+}$. Fix $\varepsilon>0$ and define $\omega_{\varepsilon}(x)=e^{-2\varepsilon|x_{1}|}$. For $R>a$ set $\Omega_{R}=\{x:a<|x|<R\}$. Consider on $H^{1}_{0}(\Omega_{R})$ the weighted form
$q_{\varepsilon}(w,v)=\int_{\Omega_{R}}\omega_{\varepsilon}\,\nabla w\cdot\nabla\overline{v}\,dx+\int_{\Omega_{R}}\omega_{\varepsilon}\,(Fx_{1}-z)w\,\overline{v}\,dx,$
which is continuous since $|\omega_{\varepsilon}Fx_{1}|\leq C_{\varepsilon}\omega_{\varepsilon}$. Adding $\lambda\int_{\Omega_{R}}\omega_{\varepsilon}w\overline{v}\,dx$ yields coercivity; hence a unique $w_{R}\in H^{1}_{0}(\Omega_{R})$ exists with boundary trace $\phi$ on $S_{a}$ and zero on $|x|=R$. Testing with $v=w_{R}$ and using $\text{{\rm Im}}~z\neq 0$ gives uniform bounds in $R$. Extending $w_{R}$ by zero outside $\Omega_{R}$ and passing to a weak limit as $R\to\infty$ (Rellich compactness on bounded sets), one obtains $u^{+}\in H^{1}_{\mathrm{loc}}(D^{+})$ with $L^{2}$ decay enforced by $\omega_{\varepsilon}$. Elliptic interior regularity implies $u^{+}\in H^{2}_{\mathrm{loc}}(D^{+})$.
Uniqueness in $D^{\pm}$ follows by taking $v=u^{\pm}$ in the weak formulation and using $\text{{\rm Im}}~z\neq 0$, while continuity with respect to $\phi$ is obtained from the uniform estimates established above. $\Box$
Lemma 4 (Surjectivity of $\Gamma_{0}$ on the defect space and definition of $\gamma_{F}(z)$).
Here $\Gamma_{0}$ and $\Gamma_{1}$ denote the boundary maps defined above, that is, $\Gamma_{0}$ is the jump of the normal derivative across $S_{a}$ and $\Gamma_{1}$ is the average trace on $S_{a}$. For $z\in\mathbb{C}\setminus\mathbb{R}$ let $N_{z}=\ker(T^{*}-z)$. Then the map
$\Gamma_{0}:N_{z}\longrightarrow H^{-1/2}(S_{a})$
is bijective. Its inverse
$\gamma_{F}(z):H^{-1/2}(S_{a})\longrightarrow N_{z}$
is bounded and satisfies $\Gamma_{0}\gamma_{F}(z)=I$. Moreover,
$M_{F}(z):=\Gamma_{1}\gamma_{F}(z):H^{-1/2}(S_{a})\longrightarrow H^{1/2}(S_{a})$
is bounded.
Proof.
Surjectivity follows from Lemma 3, by combining the interior and exterior solutions with a common trace $\phi$. If $u\in N_{z}$ satisfies $\Gamma_{0}u=0$, then $u^{\pm}$ solve homogeneous Dirichlet problems with nonreal energy, which implies $u^{\pm}\equiv 0$. Hence $\Gamma_{0}$ is also injective. The boundedness of the inverse map follows from the closed graph theorem. $\Box$
Lemma 4 shows that $\Gamma_{0}:N_{z}\to H^{-1/2}(S_{a})$ is bijective for every nonreal $z$. Since $H^{1/2}(S_{a})$ is infinite dimensional, it follows that the deficiency indices of $T$ are infinite, so $T$ is not essentially self–adjoint. Thus self–adjoint realizations arise precisely from boundary conditions relating $\Gamma_{0}$ and $\Gamma_{1}$. In particular, when the wall strength varies along the surface, $\alpha(\omega)\in L^{\infty}(S_{a};\mathbb{R})$, we introduce the boundary multiplication operator
$\widehat{A}_{\alpha}:H^{1/2}(S_{a})\to H^{-1/2}(S_{a}),\qquad\widehat{A}_{\alpha}\phi:=\alpha\phi,$
interpreted as multiplication by $\alpha$ followed by the canonical embedding $L^{2}(S_{a})\hookrightarrow H^{-1/2}(S_{a})$. In the boundary triple framework, self–adjoint extensions are obtained by imposing the condition
$\Gamma_{0}=\widehat{A}_{\alpha}\Gamma_{1},$
which enforces the jump of the normal derivative to equal the boundary value weighted by $\alpha(\omega)$. Passing to the $L^{2}(S_{a})$ pivot we then set
$A_{\alpha}^{\flat}:=\Lambda^{-1}\widehat{A}_{\alpha}\Lambda,\qquad M_{F}^{\flat}(z):=\Lambda^{-1}M_{F}(z)\Lambda^{-1},$
and use these throughout.
3 Krein resolvent formula and resolvent estimates
Before stating the Krein resolvent formula, we record concrete expressions for the Poisson operator and the boundary Weyl function.
Lemma 5 (Concrete representation of $\gamma_{F}(z)$ and $M_{F}(z)$).
Let $\tau:H^{1}(\mathbb{R}^{d})\to H^{1/2}(S_{a})$ be the trace operator, and $\tau^{*}:H^{-1/2}(S_{a})\to H^{-1}(\mathbb{R}^{d})$ its adjoint. For $z\in\mathbb{C}\setminus\mathbb{R}$, denoting by $R_{F,0}(z)=(H_{F,0}-z)^{-1}$ the free Stark resolvent, one has
$\gamma_{F}(z)$ $=$ $R_{F,0}(z)\,\tau^{*}:H^{-1/2}(S_{a})\to L^{2}(\mathbb{R}^{d}),$
$M_{F}(z)$ $=$ $\tau\,R_{F,0}(z)\,\tau^{*}:H^{-1/2}(S_{a})\to H^{1/2}(S_{a}).$
In particular, $\Gamma_{0}\gamma_{F}(z)=I$ and $\Gamma_{1}\gamma_{F}(z)=M_{F}(z)$. Moreover, $\gamma_{F}(z)$ is compact as a map $H^{-1/2}(S_{a})\to L^{2}(\mathbb{R}^{d})$.
With the operators $\gamma_{F}(z)$ and $M_{F}(z)$ from Lemma 5, we can state the Krein resolvent formula.
Theorem 6 (Krein–type resolvent formula).
Let
$A_{0}:=T^{*}|_{\ker\Gamma_{0}},\qquad A_{\alpha}:=T^{*}|_{\{u\in D(T^{*}):\Gamma_{0}u=A_{\alpha}^{b}\Gamma_{1}u\}}.$
Then $(H^{-1/2}(S_{a}),\Gamma_{0},\Gamma_{1})$ is a quasi boundary triple for $T^{*}$, and $A_{0}$, $A_{\alpha}$ are self–adjoint. Moreover, for $z\in\rho(A_{0})\cap\rho(A_{\alpha})$,
$(A_{\alpha}-z)^{-1}=(A_{0}-z)^{-1}+\gamma_{F}(z)\,\Lambda\bigl{(}I-A_{\alpha}^{\flat}M_{F}^{\flat}(z)\bigr{)}^{-1}A_{\alpha}^{\flat}\,\Lambda^{-1}\,\gamma_{F}(\bar{z})^{*}.$
In our setting, $A_{0}=H_{F,0}$, and $A_{\alpha}=H_{F,\alpha}$. Passing to the $L^{2}(S_{a})$ pivot with $\Lambda=(I-\Delta_{S_{a}})^{1/4}$ and
$A_{\alpha}^{\flat}:=\Lambda^{-1}A_{\alpha}^{b}\Lambda,\qquad M_{F}^{\flat}(z):=\Lambda^{-1}M_{F}(z)\Lambda^{-1},$
one obtains the Krein resolvent formula in the $L^{2}$ setting as above.
Remark 1.
(i) The construction of the quasi boundary triple and the Krein resolvent formula does not rely on the spherical symmetry of $S_{a}$, but only on smoothness and compactness of the boundary. Hence the results extend to $\delta$–interactions supported on smooth compact hypersurfaces [14].
(ii) The Krein resolvent formula expresses the resolvent of $H_{F,\alpha}$ as the free resolvent plus a boundary correction, so the wall appears only through operators on $L^{2}(S_{a})$.
Theorem 7 (Weighted resolvent bounds and LAP for $H_{F,0}$).
Let $d\geq 2$. For every $s>1/2$ and every $z\in\mathbb{C}\setminus\mathbb{R}$, the operator $\langle x\rangle^{-s}R_{F,0}(z)\langle x\rangle^{-s}$ is bounded on $L^{2}(\mathbb{R}^{d})$. Moreover, for $\lambda\in\mathbb{R}$ the boundary values $R_{F,0}(\lambda\pm i0)$ exist as bounded operators $\langle x\rangle^{-s}L^{2}\to\langle x\rangle^{s}L^{2}$ and depend locally Hölder–continuously on $\lambda$.
This result is classical for Stark Hamiltonians; see Herbst [12, 13] and Bentosela–Carmona–Duclos–Simon–Souillard–Weder [4].
Lemma 8 (Compactness of the boundary coupling).
Let $\text{{\rm Im}}z\neq 0$. For any $s>1/2$, the operator $\langle x\rangle^{-s}R_{F,0}(z)\langle x\rangle^{-s}$ is bounded on $L^{2}(\mathbb{R}^{d})$, and $\langle x\rangle^{s}\tau^{*}$ has compact support near $S_{a}$. Hence
$\gamma_{F}(z)=\langle x\rangle^{-s}R_{F,0}(z)\langle x\rangle^{-s}\,\langle x\rangle^{s}\tau^{*}$
is compact from $H^{-1/2}(S_{a})$ to $L^{2}(\mathbb{R}^{d})$. In particular,
$(H_{F,\alpha}-z)^{-1}-(H_{F,0}-z)^{-1}$
is compact.
Proof.
The factorization follows from Lemma 5. The boundedness of
$\langle x\rangle^{-s}R_{F,0}(z)\langle x\rangle^{-s}$
is Theorem 7. The map $\langle x\rangle^{s}\tau^{*}:H^{-1/2}(S_{a})\to H^{-1}(\mathbb{R}^{d})$ has compact support near $S_{a}$. The resolvent $R_{F,0}(z)$ maps $H^{-1}_{\mathrm{comp}}$ into $H^{1}_{\mathrm{loc}}$, and on bounded sets the embedding $H^{1}\hookrightarrow L^{2}$ is compact. Combining these maps shows that $\gamma_{F}(z)$ is compact. Finally, the Krein formula expresses the resolvent difference as a product of a compact operator, a bounded operator $\bigl{(}I-A_{\alpha}^{\flat}M_{F}^{\flat}(z)\bigr{)}^{-1}$ on $L^{2}(S_{a})$, and another bounded operator, hence it is compact. $\Box$
Remark 2.
From the physical point of view, the non–self–adjointness of $T$ reflects the fact that additional boundary conditions at the wall are required to specify the dynamics. The self–adjoint extensions $H_{F,\alpha}$ implement these conditions by prescribing a linear relation between the average boundary values and the normal jumps, which models the penetrable wall of variable strength $\alpha(\omega)$.
4 Spectrum of the penetrable–wall Stark Hamiltonian
For the case $F=0$, the following classical result is known.
Theorem 9 (Ikebe–Shimada [14]).
Let
$H_{0,\alpha}=-\Delta+A_{\alpha}\delta(|x|-a),$
acting in $L^{2}(\mathbb{R}^{3})$. Then the following hold:
  • (i)
    The essential spectrum is
    $\sigma_{\mathrm{ess}}(H_{0,\alpha})=[0,\infty).$
  • (ii)
    The negative spectrum consists of finitely many eigenvalues of finite multiplicity, which may accumulate only at 0.
  • (iii)
    The nonnegative halfline $0,\infty)$ belongs to the absolutely continuous spectrum.
Theorem 10.
Spectrum of $H_{F,\alpha}$. Let $\alpha\in L^{\infty}(S_{a};\mathbb{R})$ and $d\geq 2$. If $F\neq 0$, then
$\sigma(H_{F,\alpha})=\mathbb{R},\qquad\sigma_{\mathrm{ac}}(H_{F,\alpha})=\mathbb{R},\qquad\sigma_{\mathrm{pp}}(H_{F,\alpha})=\sigma_{\mathrm{sc}}(H_{F,\alpha})=\emptyset.$
Proof.
The argument proceeds in three steps.
Step 1: Essential spectrum. By Lemma 8 the resolvent difference $(H_{F,\alpha}-z)^{-1}-(H_{F,0}-z)^{-1}$ is compact, and hence Weyl’s stability theorem yields
$\sigma_{\mathrm{ess}}(H_{F,\alpha})=\sigma_{\mathrm{ess}}(H_{F,0}).$
Since $\sigma(H_{F,0})=\mathbb{R}$ (Herbst [12, 13]), we conclude that $\sigma_{\mathrm{ess}}(H_{F,\alpha})=\mathbb{R}$.
Step 2: Functional calculus using almost analytic extensions.
Let $\varphi\in C_{0}^{\infty}(\mathbb{R})$ and choose an almost analytic extension $\widetilde{\varphi}\in C_{0}^{\infty}(\mathbb{C})$ supported in a compact set $\Omega\Subset\mathbb{C}$ with $\widetilde{\varphi}|_{\mathbb{R}}=\varphi$ and
$|\bar{\partial}\widetilde{\varphi}(x+iy)|\leq C_{N}\,|y|^{N}\qquad(x+iy\in\mathbb{C})$ (4.1)
for some $N\geq 2$. A standard construction is
$\widetilde{\varphi}(x+iy)=\sum_{k=0}^{N-1}\frac{(iy)^{k}}{k!}\,\varphi^{(k)}(x)\,\chi(y)\,\eta(x),$
with $\chi,\eta\in C_{0}^{\infty}(\mathbb{R})$ equal to $1$ near 0 and on a neighborhood of $\mathrm{supp}\,\varphi$, respectively; then (4.1) holds.
By the Helffer–Sjöstrand formula,
$\varphi(H)=\frac{1}{\pi}\int_{\mathbb{C}}\bar{\partial}\widetilde{\varphi}(z)\,(H-z)^{-1}\,dx\,dy,\qquad\varphi(H_{0})=\frac{1}{\pi}\int_{\mathbb{C}}\bar{\partial}\widetilde{\varphi}(z)\,(H_{0}-z)^{-1}\,dx\,dy,$ (4.2)
with integrals converging in operator norm. Subtracting these identities gives
$\varphi(H)-\varphi(H_{0})=\frac{1}{\pi}\int_{\mathbb{C}}\bar{\partial}\widetilde{\varphi}(z)\,\big{(}(H-z)^{-1}-(H_{0}-z)^{-1}\big{)}\,dx\,dy.$ (4.3)
Fix a nonreal open set $\Omega_{0}\Subset\mathbb{C}\setminus\mathbb{R}$ on which $(H-z)^{-1}-(H_{0}-z)^{-1}$ is compact (by Lemma 8), and choose $\widetilde{\varphi}$ so that $\mathrm{supp}\,\bar{\partial}\widetilde{\varphi}\subset\Omega_{0}$. Then the integrand in (4.3) is compact for every $z$ in the domain of integration. Moreover, using $\|(H-z)^{-1}\|\leq|\text{{\rm Im}}~z|^{-1}$ and (4.1) we get
$\big{\|}\bar{\partial}\widetilde{\varphi}(z)\big{(}(H-z)^{-1}-(H_{0}-z)^{-1}\big{)}\big{\|}\;\leq\;C_{N}\,|\,\text{{\rm Im}}~z\,|^{N-2},$
which is integrable on $\Omega_{0}$ as soon as $N\geq 2$. Hence the Bochner integral (4.3) is a compact operator, and we have shown:
$\varphi(H)-\varphi(H_{0})\ \text{is compact for all }\ \varphi\in C_{0}^{\infty}(\mathbb{R}).$ (4.4)
This application of the Helffer–Sjöstrand functional calculus is standard; see Davies [7]. Together with the limiting absorption principle for the free Stark operator and general results of Mourre theory (Mourre [17], Amrein–Boutet de Monvel–Georgescu [1], and Jensen–Mourre–Perry [15]), the compactness property above allows one to transfer the strict Mourre estimate from $H_{F,0}$ to $H_{F,\alpha}$ up to a compact remainder.
To pass to spectral projections, let $J\subset\mathbb{R}$ be a bounded interval and choose $\{\varphi_{\varepsilon}\}_{\varepsilon\downarrow 0}\subset C_{0}^{\infty}(\mathbb{R})$ such that $0\leq\varphi_{\varepsilon}\leq 1$, $\varphi_{\varepsilon}\to\chi_{J}$ pointwise, and the supports of $\bar{\partial}\widetilde{\varphi}_{\varepsilon}$ are contained in the same fixed compact set $\Omega_{0}$, with bounds (4.1) uniform in $\varepsilon$. Then (4.3) and the dominated convergence implied by (4.1) yield
$\|\varphi_{\varepsilon}(H)-\varphi_{\varepsilon}(H_{0})\ -\ (E_{J}(H)-E_{J}(H_{0}))\|\ \longrightarrow\ 0,$
so $E_{J}(H)-E_{J}(H_{0})$ is the norm–limit of compact operators and hence compact. In particular, for later use we shall apply this with $H=H_{F,\alpha}$ and $H_{0}=H_{F,0}$.
Step 3: Mourre estimate and absence of eigenvalues. Take $A_{F}:=\mathrm{sgn}(F)\,p_{1}$ with $p_{1}=-i\partial_{x_{1}}$. Then $\Delta,\,iA_{F}]=0$ and $Fx_{1},\,iA_{F}]=|F|$, hence $H_{F,0},\,iA_{F}]=|F|I$. By [15, Def. 2.1] the pair $(H_{F,\alpha},A_{F})$ is of class $C^{1}$ on any bounded interval: indeed $H_{F,0},iA_{F}]=|F|$ and the boundary perturbation $A_{\alpha}\delta(|x|-a)$ is $H_{F,0}$–compact by Lemma 8, so the commutator with $A_{F}$ is well defined in the form sense after spectral localization. Together with the compactness of $E_{J}(H_{F,\alpha})-E_{J}(H_{F,0})$ (Step 2), the abstract Mourre theory [15, Thm. 2.2 & Thm. 4.3] (see also [1, Chap. 7]) yields the transfer of the strict Mourre estimate from $H_{F,0}$ to $H_{F,\alpha}$ up to a compact remainder. In particular,
$E_{J}(H_{F,\alpha})[H_{F,\alpha},iA_{F}]E_{J}(H_{F,\alpha})=|F|\,E_{J}(H_{F,\alpha})+R_{J},\qquad\|R_{J}\|<|F|/2,$
where $R_{J}$ is compact (the bound holds after choosing $J$ sufficiently small). Here the contribution of the boundary interaction $A_{\alpha}\delta(|x|-a)$ yields only compact terms after localization. $\Box$
5 Resonances
For $F\neq 0$ there are no eigenvalues, but former bound states at $F=0$ persist as resonances. We define resonances by applying complex distortion along the field direction and reduce the problem to the boundary.
5.1 Complex distortion and analytic continuation
The method of complex distortion (or complex scaling) provides a standard tool to define resonances in Stark-type Hamiltonians. Fix $\theta_{0}>0$ small. For $\theta\in\mathbb{C}$ with $|\text{{\rm Im}}~\theta|<\theta_{0}$ we define for $u\in C_{0}^{\infty}(\mathbb{R}^{d})$
$(U_{\theta}u)(x_{1},x_{\perp})=e^{\theta/2}\,u(e^{\theta}x_{1},x_{\perp}),\qquad x_{\perp}\in\mathbb{R}^{d-1}.$ (5.1)
For real $\theta$, $U_{\theta}$ is unitary on $L^{2}(\mathbb{R}^{d})$ and preserves $H^{2}(\mathbb{R}^{d})$. For complex $\theta$ in the strip $|\text{{\rm Im}}~\theta|<\theta_{0}$, the map $U_{\theta}$ extends by density to a boundedly invertible operator on $L^{2}$. We then set
$H_{F,0}(\theta)=U_{\theta}H_{F,0}U_{\theta}^{-1},\qquad D(H_{F,0}(\theta))=H^{2}(\mathbb{R}^{d}).$
This defines a type–A analytic family of operators in the sense of Kato. The distorted resolvent
$R_{F,0}(\theta;z)=(H_{F,0}(\theta)-z)^{-1}$
is analytic in $z$ on the resolvent set of $H_{F,0}(\theta)$.
Theorem 11 (Herbst [12, 13]).
There exists $\theta_{0}\in(0,\pi/3)$ such that the family $\theta\mapsto H_{F,0}(\theta)$ is analytic for $|\text{{\rm Im}}~\theta|<\theta_{0}$. Fix $0<\text{{\rm Im}}~\theta<\theta_{0}$. Then for every $\varepsilon\in(0,\pi/2)$ there exists $R_{\varepsilon}>0$ such that the spectrum of $H_{F,0}(\theta)$ in the angular sector
$S_{\varepsilon}\ :=\ \{\,z\in\mathbb{C}:\ -\pi+\varepsilon<\arg z<-\varepsilon,\ |z|>R_{\varepsilon}\,\}$
consists only of isolated eigenvalues of finite algebraic multiplicity. Moreover, for $f,g\in\mathcal{S}(\mathbb{R}^{d})$ the matrix elements $\langle f,(H_{F,0}-z)^{-1}g\rangle$ admit a meromorphic continuation from $\text{{\rm Im}}z>0$ to the sector $S_{\varepsilon}$, since
$\langle f,(H_{F,0}-z)^{-1}g\rangle=\langle U_{\theta}f,(H_{F,0}(\theta)-z)^{-1}U_{\theta}g\rangle.$
The poles of this continuation, equivalently the discrete eigenvalues of $H_{F,0}(\theta)$ in $S_{\varepsilon}$, are called resonances of the free Stark Hamiltonian.
Thus resonances of $H_{F,0}$ appear as eigenvalues of the distorted operator $H_{F,0}(\theta)$ lying in the lower half-plane. They are independent of the choice of distortion angle $\theta$ within the admissible strip and provide a mathematically rigorous notion of metastable states under a constant field. In the next subsection we extend this construction to the perturbed operator $H_{F,\alpha}$ by combining complex distortion with the boundary operator framework developed in Section 2.
5.2 Distorted boundary operators and Krein formula
We now extend the boundary operator framework of Section 2 to the distorted setting. Let $(\Gamma_{0},\Gamma_{1})$ denote the boundary maps introduced earlier, and let $\tau:H^{1}(\mathbb{R}^{d})\to H^{1/2}(S_{a})$ be the trace operator, $\tau^{*}:H^{-1/2}(S_{a})\to H^{-1}(\mathbb{R}^{d})$ its adjoint. For $z\in\mathbb{C}\setminus\mathbb{R}$ and $0<\text{{\rm Im}}\theta<\theta_{0}$, define the distorted Poisson operator and boundary Weyl function by
$\gamma_{F}(\theta;z):=R_{F,0}(\theta;z)\,\tau^{*},\qquad M_{F}(\theta;z):=\tau\,R_{F,0}(\theta;z)\,\tau^{*},$
where $R_{F,0}(\theta;z)=(H_{F,0}(\theta)-z)^{-1}$ is the distorted free resolvent. As before, it is convenient to work on the $L^{2}(S_{a})$ pivot using
$M_{F}^{\flat}(\theta;z):=\Lambda^{-1}M_{F}(\theta;z)\Lambda^{-1},\qquad\Lambda=(I-\Delta_{S_{a}})^{1/4}:L^{2}(S_{a})\to H^{1/2}(S_{a}).$
The distorted extension corresponding to the penetrable wall is defined by the boundary condition
$\Gamma_{0}u=A_{\alpha}^{b}\,\Gamma_{1}u,$
and is denoted by $H_{F,\alpha}(\theta)$. For $0<\text{{\rm Im}}\theta<\theta_{0}$, $U_{\theta}$ is boundedly invertible on $L^{2}$ and the distorted free operator $H_{F,0}(\theta)=U_{\theta}H_{F,0}U_{\theta}^{-1}$ is mm–sectorial [12, 13]. We keep the real boundary maps $(\Gamma_{0},\Gamma_{1})$ from Section 2, so the $\delta$–shell boundary condition $\Gamma_{0}=A_{\alpha}^{b}\Gamma_{1}$ remains unchanged under the distortion. What changes is only the Poisson and Weyl operators $\gamma_{F}(\theta;z)$ and $M_{F}(\theta;z)$, which are now constructed using the distorted free resolvent $R_{F,0}(\theta;z)$. Within the quasi boundary triple framework this leads to a closed mm–sectorial extension, whose resolvent is given by the distorted Kreĭn formula (see (5.2)); cf. [3, Sec. 6, Thm. 6.16] and [2, Thm. 2.6]. For $\theta\in\mathbb{R}$ we recover the undistorted self–adjoint extension $H_{F,\alpha}$.
Lemma 12 (Green’s identity after distortion).
Let $0<\text{{\rm Im}}\theta<\theta_{0}$ and let $T(\theta):=U_{\theta}TU_{\theta}^{-1}$. Then the adjoint $T(\theta)^{*}$ together with the boundary maps $(\Gamma_{0},\Gamma_{1})$ from Section 2 satisfies the abstract Green’s identity
$(T(\theta)^{*}u,v)-(u,T(\theta)^{*}v)=\langle\Gamma_{1}u,\Gamma_{0}v\rangle_{S_{a}}-\langle\Gamma_{0}u,\Gamma_{1}v\rangle_{S_{a}},\qquad u,v\in D(T(\theta)^{*}).$
Theorem 13 (Distorted Krein resolvent formula).
Let $A_{\alpha}^{b}:H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ be the boundary multiplication by $\alpha$ and define
$A_{\alpha}^{\flat}:=\Lambda^{-1}A_{\alpha}^{b}\Lambda,\qquad M_{F}^{\flat}(\theta;z):=\Lambda^{-1}M_{F}(\theta;z)\Lambda^{-1}.$
For $z\in\rho(H_{F,0}(\theta))\cap\rho(H_{F,\alpha}(\theta))$ one has
$(H_{F,\alpha}(\theta)-z)^{-1}=R_{F,0}(\theta;z)\;+\;\gamma_{F}(\theta;z)\,\Lambda\,\bigl{(}I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z)\bigr{)}^{-1}A_{\alpha}^{\flat}\,\Lambda^{-1}\,\gamma_{F}(\theta;\bar{z})^{*}.$ (5.2)
Proof of Theorem 13.
Fix $0<\text{{\rm Im}}\theta<\theta_{0}$ and set $T(\theta):=U_{\theta}TU_{\theta}^{-1}$. Retain the boundary maps $(\Gamma_{0},\Gamma_{1})$ of Section 2. For $z\in\mathbb{C}\setminus\mathbb{R}$ recall
$\gamma_{F}(\theta;z)=R_{F,0}(\theta;z)\tau^{*},\qquad M_{F}(\theta;z)=\tau R_{F,0}(\theta;z)\tau^{*},$
and let
$A_{0}(\theta):=T(\theta)^{*}{\upharpoonright\ker\Gamma_{0}},\qquad A_{\alpha}(\theta):=T(\theta)^{*}{\upharpoonright\{u\in D(T(\theta)^{*}):\Gamma_{0}u=A_{\alpha}^{b}\Gamma_{1}u\}}.$
Step 1 (Basic identities). By construction, $(H_{F,0}(\theta)-z)\gamma_{F}(\theta;z)\varphi=\tau^{*}\varphi$ in $\mathcal{D}^{\prime}(\mathbb{R}^{d})$, hence $\Gamma_{0}\gamma_{F}(\theta;z)=I$ and $\Gamma_{1}\gamma_{F}(\theta;z)=M_{F}(\theta;z)$.
Step 2 (Adjoint identity). For $u=\gamma_{F}(\theta;z)\varphi$ and $v\in D(A_{0}(\theta))=\ker\Gamma_{0}$, Green’s identity gives
$(\gamma_{F}(\theta;z)\varphi,f)=-\langle\varphi,\Gamma_{1}(A_{0}(\theta)-\bar{z})^{-1}f\rangle_{S_{a}},\qquad f\in L^{2}(\mathbb{R}^{d}),$
so that
$\gamma_{F}(\theta;z)^{*}=-\Gamma_{1}(A_{0}(\theta)-\bar{z})^{-1}.$
Step 3 (Ansatz and boundary equation). Let $f\in L^{2}(\mathbb{R}^{d})$ and $z\in\rho(A_{0}(\theta))\cap\rho(A_{\alpha}(\theta))$. Set $u_{0}=(A_{0}(\theta)-z)^{-1}f$, so $(T(\theta)^{*}-z)u_{0}=f$ and $\Gamma_{0}u_{0}=0$. Seek $u=u_{0}-\gamma_{F}(\theta;z)\varphi$ with unknown $\varphi\in H^{-1/2}(S_{a})$. Then $\Gamma_{0}u=-\varphi$, $\Gamma_{1}u=\Gamma_{1}u_{0}-M_{F}(\theta;z)\varphi$. Imposing the boundary condition $\Gamma_{0}=A_{\alpha}^{b}\Gamma_{1}$ yields
$\Gamma_{1}u_{0}=(M_{F}(\theta;z)-A_{\alpha}^{b})\varphi,$
hence (since $z\in\rho(A_{\alpha}(\theta))$)
$\varphi=-(I-A_{\alpha}^{b}M_{F}(\theta;z))^{-1}\Gamma_{1}u_{0}.$
Step 4 (Passage to the $L^{2}$ pivot). Note the similarity identity
$\Lambda^{-1}(I-A_{\alpha}^{b}M_{F}(\theta;z))\Lambda=I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z),$
whence
$(I-A_{\alpha}^{b}M_{F}(\theta;z))^{-1}=\Lambda(I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z))^{-1}\Lambda^{-1}.$
Inserting this into the ansatz for $u$ and recalling that $u=(H_{F,\alpha}(\theta)-z)^{-1}f$ yields (5.2), which completes the proof. $\Box$
Thus, the Krein formula continues to hold after complex distortion, providing a link between the distorted resolvent of the perturbed operator and that of the free Stark Hamiltonian. This representation will be the starting point for the boundary characterization of resonances in the next subsection.
5.3 Boundary characterization of resonances
We now characterize resonances of $H_{F,\alpha}$ through the boundary operators introduced above. Throughout we fix $0<\text{{\rm Im}}~\theta<\theta_{0}$ and work with the distorted objects $\gamma_{F}(\theta;z)$, $M_{F}(\theta;z)$ and their $L^{2}(S_{a})$ reduction $M_{F}^{\flat}(\theta;z)=\Lambda^{-1}M_{F}(\theta;z)\Lambda$.
Theorem 14 (Boundary characterization of resonances).
Fix $0<\text{{\rm Im}}\theta<\theta_{0}$ and $\text{{\rm Im}}z_{0}<0$. Then the following are equivalent:
  • (i)
    $z_{0}\in\sigma_{\mathrm{p}}(H_{F,\alpha}(\theta))$.
  • (ii)
    There exists $0\neq\varphi\in H^{-1/2}(S_{a})$ such that
    $(I-A_{\alpha}M_{F}(\theta;z_{0}))\,\varphi=0,$
    equivalently $\ker(I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z_{0}))\neq\{0\}$ in $L^{2}(S_{a})$.
  • (iii)
    For any p>d1p>d-1,
    $D_{F}^{(p)}(\theta;z):=\det\nolimits_{p}\!\bigl{(}I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z)\bigr{)}$
    is analytic on $\mathbb{C}\setminus\mathbb{R}$ and $D_{F}^{(p)}(\theta;z_{0})=0$.
Proof.
[…]
$\Gamma_{0}u=\widehat{A}_{\alpha}M_{F}(\theta;z_{0})\Gamma_{0}u\quad\Longleftrightarrow\quad\bigl{(}I-\widehat{A}_{\alpha}M_{F}(\theta;z_{0})\bigr{)}\Gamma_{0}u=0.$
Setting $\varphi:=\Gamma_{0}u\neq 0$ gives (ii).
[…]
$K(\theta;z):=A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z)\in\mathfrak{S}_{p},\qquad p>d-1,$
so $I-K(\theta;z)$ is a Fredholm operator of index zero. Since $z\mapsto M_{F}^{\flat}(\theta;z)$ is analytic on $\mathbb{C}\setminus\mathbb{R}$, the determinant
$D_{F}^{(p)}(\theta;z):=\det\nolimits_{p}\!\bigl{(}I-K(\theta;z)\bigr{)}$
is analytic as well. Standard properties of $\det_{p}$ imply
$\ker\bigl{(}I-K(\theta;z_{0})\bigr{)}\neq\{0\}\ \Longleftrightarrow\ D_{F}^{(p)}(\theta;z_{0})=0.$
Since condition (ii) in $L^{2}$–pivot form is precisely $\ker\bigl{(}I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z_{0})\bigr{)}\neq\{0\}$, this is equivalent to (iii). $\Box$
Remark 3 (Schatten class and analyticity).
For $z\notin\mathbb{R}$, the boundary operator $M_{F}^{\flat}(\theta;z)$ is compact on $L^{2}(S_{a})$; in fact, it is a classical pseudodifferential operator of order 1-1 on the smooth compact manifold $S_{a}$. Hence $M_{F}^{\flat}(\theta;z)\in\mathfrak{S}_{p}$ for all $p>2$ (and in particular $p=3$ when $d=3$), but in general not in $\mathfrak{S}_{2}$. Therefore the appropriate regularized determinant is
$\det\nolimits_{p}\!\bigl{(}I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z)\bigr{)}$
with some $p>2$ (for example $\det\nolimits_{3}$ in three dimensions). The technical background for this definition is given in Simon [19]. Boundary interaction models and the membership in Schatten classes are discussed in detail in Behrndt–Langer–Lotoreichik [2], while classical estimates on singular numbers can be found in Birman–Solomyak [5].
Remark 4 (Independence of the distortion angle).
By the Herbst distortion theory (Theorem in the previous subsection), if $0<\text{{\rm Im}}~\theta<\theta_{0}$, the discrete eigenvalues of $H_{F,\alpha}(\theta)$ in the lower half-plane are independent of $\theta$. Therefore, the set of resonances of $H_{F,\alpha}$ is well defined and coincides with the zero set of $D_{F}^{(p)}(\theta;z)$ for any admissible $\theta$.
5.4 Small–field asymptotics in three dimensions
We finally examine the fate of bound states under the application of a small Stark field. Throughout we assume $d=3$.
While the general boundary operator framework applies to $\delta$–interactions on smooth compact hypersurfaces, in this section we specialize to the case of a spherical shell. This restriction allows us to exploit spherical symmetry and obtain explicit formulas for bound states and their continuation to resonances.
Step 1: Setup at $F=0$. Suppose $E_{0}<0$ is a simple eigenvalue of $H_{0,\alpha}$ with boundary eigenvector $\phi_{0}^{\flat}\in L^{2}(S_{a})$ normalized by
$\bigl{(}(M_{0}^{\flat})^{\prime}(E_{0})\phi_{0}^{\flat},\phi_{0}^{\flat}\bigr{)}_{L^{2}(S_{a})}=1,\quad M_{0}^{\flat}(z)=\Lambda^{-1}\tau R_{0}(z)\tau^{*}\Lambda^{-1},\quad R_{0}(z)=(-\Delta-z)^{-1}.$ (5.3)
Step 2: Analytic expansion of $M_{F}(z)$. For $|F|$ sufficiently small, the resolvent admits the Neumann expansion
$R_{F,0}(z)=(H_{0}+Fx_{1}-z)^{-1}=R_{0}(z)-FR_{0}(z)\,x_{1}\,R_{0}(z)+F^{2}R_{0}(z)\,x_{1}R_{0}(z)\,x_{1}R_{0}(z)$-$\cdots$
valid for $z\in\mathbb{C}\setminus[0,\infty)$. Substituting into $M_{F}(z)=\tau R_{F,0}(z)\tau^{*}$ gives
$M_{F}(z)$ $=$ $M_{0}(z)+FM_{1}(z)+F^{2}M_{2}(z)+O(F^{3}),$
$M_{1}(z)$ $=$ $-\tau R_{0}(z)\,x_{1}\,R_{0}(z)\tau^{*},$
$M_{2}(z)$ $=$ $\tau R_{0}(z)\,x_{1}R_{0}(z)\,x_{1}R_{0}(z)\tau^{*}.$
Step 3: Analytic perturbation of the eigenvalue. Let $G(F,z)=I+A_{\alpha}M_{F}(z)$. Since $E_{0}$ is a simple zero of $\det(\Lambda^{-1}G(0,z)\Lambda)$, the implicit function theorem yields an analytic branch $z(F)=E_{0}+Fa_{1}+F^{2}a_{2}+O(F^{3})$ with
$a_{1}=-\bigl{(}M_{1}^{\flat}(E_{0})\phi_{0}^{\flat},\phi_{0}^{\flat}\bigr{)}_{L^{2}(S_{a})}.$ (5.4)
Step 4: Exponentially small resonance widths. Let $\phi_{0}^{\flat}\in L^{2}(S_{a})$ be the normalized boundary eigenvector corresponding to the simple eigenvalue $E_{0}<0$ of $H_{0,\alpha}$ as in Step 1. Then the branch $z(F)$ obtained in Step 3 satisfies $\text{{\rm Re}}z(F)\to E_{0}$ as $F\downarrow 0$. It is known from semiclassical tunneling theory that the imaginary part $\text{{\rm Im}}z(F)$ is exponentially small in $1/F$. To make this more explicit, we recall that the exponential factor can be computed in terms of an Agmon–type action integral along the field direction.
Agmon distance along the field (derivation).
In the exterior region $\{x_{1}>a\}$ the effective potential is $V(x_{1})=Fx_{1}$, and for the bound state energy $E_{0}<0$ (at $F=0$) the Agmon action along the straight path in the field direction is
$d(a\to x_{1})=\int_{a}^{x_{1}}\!\sqrt{Fs-E_{0}}\,ds=\frac{2}{3F}\Big{[}(Fx_{1}-E_{0})^{3/2}-(Fa-E_{0})^{3/2}\Big{]}.$
Thus the exterior WKB tail behaves like $|u(x_{1})|\propto\exp(-d(a\to x_{1}))$, and since $d(a\to x_{1})\sim\mathrm{const}/F$ as $F\downarrow 0$, the leakage across the barrier —and hence the resonance width— is exponentially small in $1/F$. This explicit formula clarifies how the exponential suppression arises from the tunneling barrier created by the linear potential. In particular, it explains why the resonance width behaves like $\exp(-c/F)$ as $F\downarrow 0$. More precisely:
Theorem 15 (Exponentially small resonance widths for small fields).
Let $d=3$ and let $E_{0}<0$ be a simple eigenvalue of $H_{0,\alpha}$. There exist constants $F_{0},c_{1},c_{2}>0$ such that for all $0<F<F_{0}$ the operator $H_{F,\alpha}$ has a resonance $z(F)$ with $\text{{\rm Re}}z(F)\to E_{0}$ as $F\downarrow 0$ and
$e^{-c_{2}/F}\ \lesssim\ \Gamma(F):=-2\,\text{{\rm Im}}~z(F)\ \lesssim\ e^{-c_{1}/F}.$
The constants $c_{1},c_{2}$ can be expressed in terms of an Agmon-type distance along the field direction associated with the effective potential barrier.
Proof.
Let $E_{0}<0$ be a simple eigenvalue of $H_{0,\alpha}$ with normalized boundary eigenvector $\phi_{0}^{\flat}\in L^{2}(S_{a})$ as in Step 1. As explained above, the branch $z(F)$ of resonances near $E_{0}$ is obtained from the equation $G(F,z)=I+A_{\alpha}M_{F}(z)$ not invertible. Analytic perturbation theory yields
$z(F)=E_{0}+Fa_{1}+F^{2}a_{2}+O(F^{3}),\qquad a_{1}=-(M_{1}^{\flat}(E_{0})\phi_{0}^{\flat},\phi_{0}^{\flat})_{L^{2}(S_{a})},$
and $\text{{\rm Re}}z(F)\to E_{0}$ as $F\downarrow 0$. The imaginary part is more delicate.
By Herbst’s complex distortion, resonances of $H_{F,\alpha}$ correspond to discrete eigenvalues of $H_{F,\alpha}(\theta)$ for $0<\text{{\rm Im}}~\theta<\theta_{0}$. The corresponding eigenfunction $u_{F}$ satisfies $(H_{F,0}+A_{\alpha}\delta(|x|-a)-z(F))u_{F}=0$. Inside the wall, $u_{F}$ remains close to the bound state of $H_{0,\alpha}$, while in the exterior region $\{x_{1}>a\}$ it decays according to the Agmon action computed above. Applying Agmon’s method then yields the upper bound
$\Gamma(F)=-2\text{{\rm Im}}~z(F)\ \lesssim\ \exp(-c_{1}/F)$
for some $c_{1}>0$. Conversely, Carleman–type lower bounds exclude faster-than-exponential decay, giving
$\Gamma(F)\ \gtrsim\ \exp(-c_{2}/F)$
for some $c_{2}>0$. Both constants $c_{1},c_{2}$ are expressed in terms of the Agmon distance across the barrier region in the field direction.
Collecting these estimates, we obtain
$e^{-c_{2}/F}\ \lesssim\ \Gamma(F)\ \lesssim\ e^{-c_{1}/F},\qquad 0<F<F_{0},$
for some $F_{0}>0$. This proves the claim. $\Box$
This proof is a concise summary of the detailed arguments given by Wang [20, 21].
This establishes that bound states at $F=0$ bifurcate into resonances whose widths vanish faster than any power of $F$ as $F\downarrow 0$, in accordance with the general theory of metastable states under static fields.
Remark 5 (Physical interpretation for constant $\alpha<0$).
In the case of a constant attractive wall strength $\alpha<0$, the negative eigenvalue $E_{0}=-\kappa^{2}$ at $F=0$ corresponds to a bound state localized near the spherical shell. Switching on a small electric field $F>0$ tilts the potential and creates an effective barrier in the field direction. The Agmon action $d(a\to x_{1})$ measures the tunneling probability through this barrier, and the resonance width $\Gamma(F)$ corresponds to the decay rate of the metastable state. Thus the mathematical result that $\Gamma(F)\sim e^{-c/F}$ admits a direct physical interpretation: the bound state leaks out by quantum tunneling, in close analogy with the mechanism of $\alpha$–decay in nuclear physics.
Comparison with the free Stark Hamiltonian
It is instructive to compare our results with the well–known case of the free Stark Hamiltonian.
  • For the free Stark Hamiltonian, the spectrum is purely absolutely continuous and no bound states are present. Its resonances are known to exist only in an abstract sense, with widths that are exponentially small in the inverse field strength, but no explicit correspondence with eigenvalues is available.
  • By contrast, for the penetrable–wall Stark Hamiltonian studied here, bound states exist at zero field, and their continuation under a nonzero field yields resonances that can be traced explicitly. Their positions admit perturbative expansions, and their widths are again exponentially small, now governed by an Agmon–type distance determined by the wall and the associated bound state.
This comparison highlights that, unlike the free Stark Hamiltonian where resonances remain abstract, the penetrable–wall model provides a concrete and perturbative description of resonance formation and decay.
Point spectrum at zero field for constant wall strength
When the wall strength is constant, $\alpha(\omega)\equiv\alpha$, the negative spectrum of $H_{0,\alpha}$ can be determined explicitly in terms of modified spherical Bessel functions. For $l=0$, bound states at energy $E=-\kappa^{2}$ with $\kappa>0$ satisfy
$\alpha=-\kappa\bigl{(}1+\coth(\kappa a)\bigr{)}.$
From this relation one sees that only finitely many negative eigenvalues exist, accumulating at zero, in agreement with Ikebe–Shimada [14]. Note also that bound states appear only for attractive wall strength $\alpha<0$; for $\alpha>0$ the spectrum is purely $0,\infty)$.
Moreover, the equation yields simple asymptotics for $\kappa$:
  • In the weak coupling regime ($\kappa a\ll 1$),
    $\kappa\approx-\alpha-\frac{1}{a},\qquad E\approx-\left(\alpha+\frac{1}{a}\right)^{2}.$
  • In the strong coupling regime ($\kappa a\gg 1$),
    $\kappa\approx-\frac{\alpha}{2},\qquad E\approx-\frac{\alpha^{2}}{4}.$
These explicit formulas clarify the structure of the point spectrum at zero field and provide the starting point for the resonance asymptotics under a nonzero Stark field.